COHERENT CONTROL OF IONIZATION FRAGMENTATION OF SO_{2}

\underline{BING XUE}, JUN HAN, AND HAI-LUNG DAI, Department of Chemistry, University of Pennsylvania, Philadelphia, PA 19104; YANG CHEN, Department of Chemical Physics, University of Sciences and Technologies of China, Hefei, Anhui 230026, PR China.

The coherent control of ionization fragment of SO_{2} is obtained by managing the ionization contribution from a highly vibrational excited state of the ground electronic state, which is coupled to a ro-vibronic state of ^{q}R(6)_{1,5} of SO_{2} C̃ state and is coherently prepared by a pump laser pulse. Carefully choosing the probe wavelength, the ion fragment from either the ”bright” or the ”dark” state of SO_{2} can be manipulated.